OpenStudy (aravindg):

1 dimensional motion problems

OpenStudy (aravindg):

A body falling freely under the action of gravity passess 2 points 40 m apart vertically in one second.Find from what height above the upper point it began to fall

OpenStudy (aravindg):

help

OpenStudy (aravindg):

.............

OpenStudy (aravindg):

...........................................

OpenStudy (nikvist):

\[\Delta s=40m\quad,\quad\Delta t=1s\]\[s_1=\frac{1}{2}gt^2\]\[s_2=s_1+\Delta s=\frac{1}{2}g(t+\Delta t)^2\]\[\Delta s=\frac{1}{2}g(t+\Delta t)^2-\frac{1}{2}gt^2=\frac{1}{2}g(2t+\Delta t)\Delta t\]\[t=\frac{1}{2}\left(\frac{2\Delta s}{g\Delta t}-\Delta t\right)=3.5s\]\[s_1=\frac{1}{2}gt^2=61.25m\quad(g=10m/s^2)\]

OpenStudy (anonymous):

x = x0 - v0t - (gt^2)/2 x-x0 = 40 40 = -v0t -(gt^2)/2 it happened in 1 second so : 40 = -v0 - 5 v0 = -45 m/s. now we need to calculate the time that it took it to reach 45 m/s from 0 v = -at -45 = -10t t= 4.5 sec now the only thing to do is find the distance that it fell in 4.5 seconds with zero speed to begin with x=(gt^2)/2 x = (10 * (4.5)^2 )/ 2 = 101.25 m so 101.25 m above the upper point